# Integrating Multimodal Data for Joint Generative Modeling of Complex Dynamics

### v2

@@ 3.1 Evaluating MTF: Comparison to other algorithmic strategies
Figure 2: DS reconstruction from moderately (a) and heavily (b-c) distorted continuous observations (Gaussian observation noise of $10\%$ and $50\%$, respectively, of the data variance) and other simultaneously provided observation modalities, sampled from a Lorenz-63 system. (a) Freely generated example trajectories from a dendPLRNN ($M=20,B=10,K=20,\tau=10$) trained with MTF jointly on Gaussian ($10\%$ noise), ordinal, and count data. (b) Same as (a) for a dendPLRNN trained by MTF on heavily distorted Gaussian ($50\%$ noise) and ordinal observations. Note that even in this case the butterfly wing structure of the Lorenz attractor was successfully captured. The maximum Lyapunov exponent ($\lambda_{\text{max}}$) furthermore confirms the dendPLRNN-generated attractor is chaotic (for the GT Lorenz system, $\lambda_{\text{max}}\approx 0.903$). (c) Normalized cumulative histograms of geometrical attractor disagreement ($D_{stsp}$, left) and Hellinger distance ($D_{H}$, right) between reconstructed and ground-truth system for the same setting as in (b).
3.2 Dynamical systems reconstruction from discrete random variables
Motivated by these results, we pushed the system even further and attempted DS reconstruction solely from ordinal data (created as above), i.e. completely omitting continuous observations. This is profoundly more challenging than the multimodal setting with Gaussian data, since the ordinal process considerably coarse-grains the underlying continuous dynamical process. Since in this case we do not have a direct linear mapping between ground truth state space and that of the trained dendPLRNN (which in the case of Gaussian observations would simply be given by the linear decoder, eq. 11), we construct one post-hoc by optimizing a linear operator given by a linear dimensionality reduction (PCA) concatenated with a geometry-preserving rotation operation (see Methods A.2). Fig. 3 (top) confirms that, using MTF, successful DS reconstruction is still possible under these conditions.
Finally, we asked whether DS reconstruction might be feasible based on a purely symbolic representation of the dynamics. For this, $4^{3}$ symbols were used corresponding to subregions defined by a $4\times 4\times 4$ grid superimposed on the attractor (this symbolic code was reduced further and then delay-embedded, see Suppl. A.5 for full details). Fig. 3 (bottom) illustrates the general procedure, and also shows that, using MTF, successful DS reconstruction is – in principle – possible from just a symbolic (categorical) coding of the underlying chaotic attractor. This is to our knowledge the first time such a result has been shown. Comparable results could not be achieved by the multimodal sequential VAE or MS (see Table B4). Fig. A8 shows that even the maximum Lyapunov exponent $\lambda_{max}$ of the system reconstructed with MTF from just the symbolic coding of the Lorenz attractor often comes close to the true value (see also Fig. A6).
@@ 3.2 Dynamical systems reconstruction from discrete random variables
Figure 3: DS reconstruction from discrete observations by MTF ($M=30,B=15,K=30,\tau=10$). Top: Reconstruction of Rössler attractor from only ordinal time series. Bottom: DS reconstruction from symbolic coding of Lorenz attractor (see Fig. A7 for true and predicted class label probabilities). Note that in both cases the topology and general geometry are preserved, and maximal Lyapunov exponents closely match those of the true systems (Rössler: $\lambda_{\text{max}}^{\text{true}}\approx 0.072$, Lorenz: $\lambda_{\text{max}}^{\text{true}}\approx 0.903$). TDE = temporal delay embedding.
3.3 Reconstruction from neuroimaging and behavioral data
@@
We evaluated MTF’s performance on real multimodal time series using a dataset of functional magnetic resonance imaging (fMRI) recordings from human subjects undertaking various cognitive tasks in an fMRI scanner (Koppe et al., 2014). The data consisted of continuous blood-oxygenation-level-dependent (BOLD) time series acquired from 26 subjects at a sampling rate of $1/3$ $\mathrm{Hz}$. The first principal component of BOLD activity from each of 20 different brain regions was selected for analysis. The individual time series were rather short ($T=360$ per subject), posing a significant challenge for DS reconstruction. In addition to continuous BOLD activity, categorical time series of cognitive stage labels were given for each subject, corresponding to the five cognitive tasks that each subject went through repeatedly during the experiment (‘Rest’, ‘Instruction’, ‘Choice Reaction Task [CRT]’, ‘Continuous Delayed Response Task [CDRT]’ and ‘Continuous Matching Task [CMT]’) [^4] Note that CRT, CDRT, and CMT did not differ in terms of presented stimuli or response types, but only in cognitive load, thus reflecting processes internal to the subject., which we accommodated via a multi-categorical decoder model, eq. 15 in Methods A.1.
[…]
@@
Fig. A2(b) further shows an example where on a left-out test set simulated BOLD activity and model-generated task labels closely resembled those of the real subject. Fig. A2(c) investigates to what extent cross-modal links have been built by the trained system in its latent space by illustrating that after multimodal training, freely sampled latent activity trajectories from the dendPLRNN reflect the structure of the different tasks stages. Fig. A10 in Suppl. A.5 provides generated time series for several example subjects. Table 2, finally, compares the DSR performance on the fMRI and behavioral data for all the methods introduced in sect. 3, with MTF substantially outperforming the sequential MVAE from (Kramer et al., 2022).
@@ 3.4 DSR from neurophysiological and location data
As a second empirical test case we used electrophysiological recordings of multiple single neurons from the CA1 region of the rodent hippocampus (Grosmark et al., 2016), a brain region involved in spatial navigation and memory (Buzsaki, 2006; Moser et al., 2017). During the experiments the rats traversed a 1.6 m long track along which the longitudinal position of the animal was recorded, treated as a continuously valued Gaussian $1$d time series. Water rewards were provided at each of the two end points of the track, indicated to the model through brief external inputs. At the same time, for the data set on which we focused our analysis, the spiking activity (action potential time points) of $120$ neurons was recorded via intracortical electrodes (of which $60$ with reasonably high spiking rate were retained). Following common procedures in neuroscience (Zhou & Wei, 2020; Schneider et al., 2023), spike trains were segmented into $200$ ms bins, resulting in a multivariate count time series (for details, see Methods A.5). These were modelled with a negative-binomial decoder (capturing the high dispersion better than Poisson or zero-inflated Poisson models).
@@
Multimodal (position and spike count) data were split into a training and a test set of $4600$ time steps each. Example true and model-generated spike trains for the test set are in Fig. 4 (a) (see Fig. A11 for training set). Fig. 4 (b) compares various standard spike train statistics like the mean firing rate, mean zero rate, coefficient of variation, and the neural cross-correlation matrix between experimental and model-generated data, iterated forward from one inferred initial state across the whole test set. The agreement in spiking statistics between model-simulated long-term behavior and test data was in fact as good as between different segments (training and test set) of the experimental data itself. Moreover, predicted position on the track closely matched that of the real behavior for the test set (Fig. 4 (a)). Using both data sources (spike trains and position) led to far better reconstructions across all spike train performance metrics, and overall more robust reconstructions, than when only spike count data were provided for model training (Mann-Whitney U-test across $50$ trained models, $p<0.025$ for all metrics; Fig. 4 (c)). In the CA1 region, place cells are known to encode spatial information (O’Keefe & Nadel, 1978). Our results indicate that the MTF algorithm can leverage this information to build a joint latent model of the neural recordings and movement of the rat (Fig. 4 (d)).
Figure 4: (a) Example reconstructions of spike trains and spatial location of animal. Top: Spike train data from a test set not used for model training (topmost), and model-generated spike trains (below) simulated from a data-inferred initial condition. Bottom: True and model-predicted spatial position on the test set. (b) Top: Correlation of various spike train statistics between test set and model-generated data (blue), and – for comparison – between experimental training and test set data (orange): mean spike rate, zero count ratio, coefficient of variation, and correlation between cross-correlation coefficients. Diagonal gray lines are bisectrices, not regression lines. Bottom: Cross-correlation matrices among all neurons for the experimental test set (left) and model-generated spike data (right). (c) Joint DSR from both spatial and neural data significantly improves reconstructions compared to just neural data alone. (d) DSR model latent space (shown is a subspace), illustrating how the latent dynamics is organized according to the animal’s spatial position (color-coded).
@@ 4 Conclusions
Here we introduced a novel training framework (MTF) for DSR from non-Gaussian and multimodal time series data. While DS reconstruction is meanwhile a large field in scientific machine learning (Brunton et al., 2016; Pathak et al., 2018; Lejarza & Baldea, 2022; Brenner et al., 2022; Hess et al., 2023; Chen & Wu, 2023), reconstruction based on discrete or multimodal data has hardly been addressed so far, although such scenarios are commonplace in many areas like medicine, neuroscience, or climate research. Here we utilize recent insights on guiding the training process by control signals, as in STF (Mikhaeil et al., 2022) and GTF (Hess et al., 2023), within a novel multimodal DSR framework. In our approach, a sparse TF signal is generated by an MVAE that translates different data modalities into a common latent code in order to guide a DSR model. Our general framework is very flexible and allows to easily swap DSR, encoder, or decoder models for the problem at hand. For diverse benchmarks and sampling conditions, MTF clearly outperforms various other training setups and models. We further highlighted on examples from neuroscience how the DSR model’s latent space exposes meaningful relations between different data modalities (neural activity and behavior in our case). In contrast to other recently proposed methods for construing a common latent space from multiple data sources (Schneider et al., 2023), our method actually delivers a generative model of the system dynamics. A crucial difference to ‘classical’ multimodal sequential VAEs (Kramer et al., 2022) is that our algorithm allows to optimally control trajectory and gradient flows whilst training, a key ingredient for successful DS reconstruction (Mikhaeil et al., 2022; Hess et al., 2023; Eisenmann et al., 2023).
@@
This way, MTF was even able, under certain conditions, to recover chaotic attractors from ordinal or categorical observations alone, including sensible estimates of their maximum Lyapunov exponent. This, to our knowledge, has never been shown before, and at first glance appears puzzling, as an ordinal or symbolic coding of the dynamics seems to remove much of the geometrical information. However, results in the field of symbolic dynamics ensure that under certain conditions a symbolic encoding of a DS preserves topological invariants and Lyapunov spectra (Hubler, 2012; Osipenko, 2000; McCullough et al., 2017; Zhang et al., 2017), and delay embedding theorems have been formulated for non-continuous signals like point processes (Sauer, 1994, 1995). While here we presented first results on DS reconstruction for discrete data, whether and how much of the original state space topology of a data-generating DS can be recovered from non-continuous, non-Gaussian random variables remains an important topic for future theoretical and empirical research.
@@ Acknowledgements
This work was funded by the European Union’s Horizon 2020 research and innovation Programme under grant agreement 945263 (IMMERSE), by the German Research Foundation (DFG) within the collaborative research center TRR-265 (project A06), by the German Research Foundation (DFG) within Germany’s Excellence Strategy EXC 2181/1 – 390900948 (STRUCTURES), and by a living lab grant by the Federal Ministry of Science, Education and Culture (MWK) of the state of Baden-Württemberg, Germany (grant number 31-7547.223-7/3/2), to DD and GK.
Ethics Statement
@@ A.1 MTF framework
The MTF framework consists of four major components, 1) a DSR model for approximating the flow of the underlying, observed DS, 2) a set of modality-specific decoder models which capture the statistical particularities of the observed time series, 3) a MVAE for integrating all different data modalities in order to produce a joint latent code which can be used for TF, and 4) the actual MTF training algorithm. Here we describe each of these components in turn.
@@
While our framework is generic and allows to work with any type of DSR model, here we used the dendPLRNN, eq. 1, recently introduced specifically for DSR (Brenner et al., 2022) (see Suppl. Table B3 for comparison with other RNN models). A major advantage of the dendPLRNN and related formulations is that it provides some degree of analytical tractability, which is of major importance in scientific settings where we are usually interested in analyzing the data-inferred model further as a formal surrogate for the real system. For example, for the dendPLRNN we have efficient algorithms for precisely locating all fixed points, cycles, and bifurcation manifolds (Eisenmann et al., 2023), and it allows for translation into an equivalent continuous-time representation (Monfared & Durstewitz, 2020), which further eases certain types of DS analysis.
@@
To infer the latent process dendPLRNN jointly from multiple data modalities, it is connected to different decoder models that take the distinct distributional properties of each modality into account, Eqn. 2-4.
@@
Categorical observations are, like ordinal observations, not associated with a metric space, and in contrast to ordinal data there is also no natural ordering between variable values. To couple categorical observations to the latent states, we employed the natural link function given by
@@
Here, the parameters $\bm{\beta}_{i}\in\mathbb{R}^{M\times 1}$ constitute the respective regression weights for category $i=1\ldots K-1$, with the total probability over all categories $\sum_{i=1}^{K}\pi_{i}=1$.
For count observations $\left\{\bm{c}_{t}\right\}_{t=1}^{T}$, with $\bm{c}_{t}=\left(c_{1t},\dots,c_{Lt}\right)^{T}$, we tried two different decoder models. First, a standard Poisson model,
@@
These probabilities are related to the latent states via the log-link function, $\log\lambda_{lt}=\gamma_{0}^{(l)}+\sum_{m=1}^{M}\gamma_{m}^{(l)}z_{mt}$, where $\bm{\gamma}^{(l)}$ is a vector of coefficients. Thus, $\lambda_{lt}=e^{\gamma_{0}^{(l)}+\bm{\gamma}^{(l)}\bm{z_{t}}}$ is the expected count for the $l^{\text{th}}$ observation at time $t$.
Second, a negative binomial model given by
@@
where $\mu_{lt}$ is the mean count and $\phi_{l}$ the dispersion parameter of the negative binomial distribution for the $l^{\text{th}}$ observation at time $t$. As for the Poisson model, we use a log-link function, $\log\mu_{lt}=\gamma_{0}^{(l)}+\sum_{m=1}^{M}\gamma_{m}^{(l)}z_{mt}$, with $\bm{\gamma}^{(l)}$ a vector of coefficients, and $z_{mt}$ the $m^{\text{th}}$ latent variable at time $t$. Properly accounting for dispersion significantly improved modeling of the spike counts. The likelihood function is based on the torch.distributions.NegativeBinomial class from the PyTorch library.

### v3

@@ 3.1 Evaluating MTF: Comparison to Other Algorithmic Strategies
Figure 2: DS reconstruction from moderately (a) and heavily (b-c) distorted continuous observations (Gaussian observation noise of $10\%$ and $50\%$, respectively, of the data variance) and other simultaneously provided observation modalities, sampled from a Lorenz-63 system. a: Freely generated example trajectories from a dendPLRNN ($M=20,B=10,K=20,\tau=10$) trained with MTF jointly on Gaussian ($10\%$ noise), ordinal, and count data. b: Same as a for a dendPLRNN trained by MTF on heavily distorted Gaussian ($50\%$ noise) and ordinal observations. Note that even in this case the butterfly wing structure of the Lorenz attractor was successfully captured. The maximum Lyapunov exponent ($\lambda_{\text{max}}$) furthermore confirms the dendPLRNN-generated attractor is chaotic (for the GT Lorenz system, $\lambda_{\text{max}}\approx 0.903$). c: Normalized cumulative histograms of geometrical attractor disagreement ($D_{stsp}$, left) and Hellinger distance ($D_{H}$, right) between reconstructed and ground-truth system for the same setting as in b.
To investigate multimodal data integration in a more challenging setting, we next tested a scenario where continuous observations from the Lorenz-63 system were heavily distorted by Gaussian noise with $50\%$ of the data variance. At the same time, ordinal observations with $8$ variables divided into $7$ ordered categories each, $o_{nt}\in\{1\dots 7\}$, $n=1\dots 8$, were sampled using Eq. 14. We then trained the dendPLRNN via MTF once with, and once without, ordinal observations. Fig. 2b proves that with ordinal observations on board, DS reconstruction is, in principle, possible even under these challenging conditions. The cumulative histograms of the geometric measure $D_{stsp}$ and the temporal measure $D_{H}$ in Fig. 2c furthermore show that inclusion of ordinal observations significantly improves reconstructions.
Figure 3: Cross-modal inference with missing observations, using a mixture-of-experts encoder. a: Reconstruction of the Lorenz-63 from Gaussian and ordinal observations, with $20\%$ of data removed at random times chosen independently for each modality. b: Using only the Gaussian expert, the corresponding ordinal observations can be decoded almost perfectly, including at times missing in the ordinal training data. Dashed lines indicate sections with missing data in a and b.
Finally, MTF can easily handle missing data in one or more modalities, by simply dropping missing data points from the loss terms and using encoders like a mixture-of-experts (Appx. A.4), as illustrated in Fig. 3. As Fig. 3 highlights as well, MTF can also be employed for cross-modal inference, predicting outcomes for unobserved modalities through the joint encoder model.
3.2 DSR from Discrete Random Variables
@@
Figure 4: DS reconstruction from discrete observations by MTF ($M=30,B=15,K=30,\tau=10$). Top: Reconstruction of Rössler attractor from only ordinal time series. Bottom: DS reconstruction from symbolic coding of Lorenz attractor (see Fig. A9 for true and predicted class label probabilities). Ground truth systems and their ordinal/symbolic encoding are on the left, corresponding reconstructions on the right. Note that in both cases the topology and general geometry are preserved, and maximal Lyapunov exponents closely match those of the true systems (Rössler: $\lambda_{\text{max}}^{\text{true}}\approx 0.072$, Lorenz: $\lambda_{\text{max}}^{\text{true}}\approx 0.903$). TDE = temporal delay embedding.
Motivated by these results, we pushed the system even further and attempted DS reconstruction solely from ordinal data (created as above), i.e. completely omitting continuous observations. This is profoundly more challenging than the multimodal setting with Gaussian data, since the ordinal process considerably coarse-grains the underlying continuous dynamical process. Since in this case we do not have a direct linear mapping between ground truth state space and that of the trained dendPLRNN (which in the case of Gaussian observations would simply be given by the linear decoder, Eq. 12), we construct one post-hoc by optimizing a linear operator given by a linear dimensionality reduction (PCA) concatenated with a geometry-preserving rotation operation (see Methods A.2). Fig. 4 (top) confirms that, using MTF, successful DS reconstruction is still possible under these conditions (see also Fig. A14 for a higher-dimensional example of reconstruction from only ordinal data).
Finally, we asked whether DS reconstruction might be feasible based on a purely symbolic representation of the dynamics. For this, $4^{3}$ symbols were used corresponding to subregions defined by a $4\times 4\times 4$ grid superimposed on the attractor (this symbolic code was reduced further and then delay-embedded, see Appx. A.5 for full details). Fig. 4 (bottom) illustrates the general procedure, and also shows that, using MTF, successful DS reconstruction is – in principle – possible from just a symbolic (categorical) coding of the underlying chaotic attractor. This is to our knowledge the first time such a result has been shown. Comparable results could not be achieved by the multimodal SVAE or MS (see Table B4). Fig. A9b shows that even the maximum Lyapunov exponent $\lambda_{max}$ of the system reconstructed with MTF from just the symbolic coding of the Lorenz attractor often comes close to the true value (see also Fig. A8).
3.3 DSR from Neuroimaging and Behavioral Data
@@
We evaluated MTF’s performance on real multimodal time series using a dataset of functional magnetic resonance imaging (fMRI) recordings from human subjects undertaking various cognitive tasks in an fMRI scanner (Koppe et al., 2014). The data consisted of continuous blood-oxygenation-level-dependent (BOLD) time series acquired from 20 subjects at a sampling rate of $1/3$ $\mathrm{Hz}$. The first principal component of BOLD activity from each of 20 different brain regions was selected for analysis. The individual time series were rather short ($T=360$ per subject), posing a significant challenge for DS reconstruction. In addition to continuous BOLD activity, categorical time series of cognitive stage labels were given for each subject, corresponding to the five cognitive tasks that each subject went through repeatedly during the experiment (‘Rest’, ‘Instruction’, ‘Choice Reaction Task [CRT]’, ‘Continuous Delayed Response Task [CDRT]’ and ‘Continuous Matching Task [CMT]’) [^4] Note that CRT, CDRT, and CMT did not differ in terms of presented stimuli or response types, but only in cognitive load, thus reflecting processes internal to the subject., which we accommodated via a multi-categorical decoder model, Eq. 16 in Methods A.1.
@@
[…]
Fig. A2b further shows an example where on a left-out test set simulated BOLD activity and model-generated task labels closely resembled those of the real subject. Fig. A2c investigates to what extent cross-modal links have been built by the trained system in its latent space by illustrating that after multimodal training, freely sampled latent activity trajectories from the dendPLRNN reflect the structure of the different tasks stages. Fig. A3 in Appx. A.5 provides generated time series for several example subjects. Table 2, finally, compares the DSR performance on the fMRI and behavioral data for all the methods introduced in Sect. 3, with MTF substantially outperforming the multimodal SVAE from (Kramer et al., 2022), again highlighting the crucial differences between these approaches.
@@ 3.4 DSR from Neurophysiological and Location Data
As a second empirical test case, we used electrophysiological recordings of multiple single neurons from the CA1 region of the rodent hippocampus (Grosmark et al., 2016), a brain region involved in spatial navigation and memory (Buzsaki, 2006; Moser et al., 2017). During the experiments the rats traversed a 1.6 m long track along which the longitudinal position of the animal was recorded, treated as a continuously valued Gaussian $1$d time series. Water rewards were provided at each of the two end points of the track, indicated to the model through brief external inputs. At the same time, for the data set on which we focused our analysis, the spiking activity (action potential time points) of $120$ neurons was recorded via intracortical electrodes (of which $60$ with reasonably high spiking rate were retained). Following common procedures in neuroscience (Zhou & Wei, 2020; Schneider et al., 2023), spike trains were segmented into $200$ ms bins, resulting in a multivariate count time series (for details, see Methods A.5). These were modeled with a negative-binomial decoder (Eq. 20), capturing the high dispersion better than Poisson or zero-inflated Poisson models.
@@
Multimodal (position and spike count) data were split into a training and a test set of $4600$ time steps each. Example true and model-generated spike trains for the test set are in Fig. 5a (see Fig. A11 for training set). Fig. 5b compares various standard spike train statistics like the mean firing rate, mean zero rate, coefficient of variation, and the neural cross-correlation matrix between experimental and model-generated data, iterated forward from one inferred initial state across the whole test set. The agreement in spiking statistics between model-simulated long-term behavior and test data was in fact as good as between different segments (training and test set) of the experimental data itself. Moreover, the predicted position on the track closely matched that of the real behavior for the test set (Fig. 5a). Using both data sources (spike trains and position) led to far better reconstructions across all spike train performance metrics, and overall more robust reconstructions, than when only spike count data were provided for model training (Mann-Whitney U-test across $50$ trained models, $p<0.025$ for all performance metrics; Fig. 5c). In the CA1 region, place cells are known to encode spatial information (O’Keefe & Nadel, 1978). Our results indicate that the MTF algorithm can leverage this information to build a joint latent model of the neural recordings and movement of the rat (Fig. 5d).
@@ 4 Conclusions
Here we introduced a novel training framework (MTF) for DSR from non-Gaussian and multimodal time series data. While DS reconstruction is meanwhile a large field in scientific machine learning (Brunton et al., 2016; Pathak et al., 2018; Lejarza & Baldea, 2022; Brenner et al., 2022; Hess et al., 2023; Chen & Wu, 2023), reconstruction based on discrete or multimodal data has hardly been addressed so far, although such scenarios are commonplace in many areas like medicine, neuroscience, or climate research. Here we utilize recent insights on guiding the training process by control signals, as in STF (Mikhaeil et al., 2022) and GTF (Hess et al., 2023), within a novel multimodal DSR framework. In our approach, a sparse TF signal is generated by an MVAE that translates different data modalities into a common latent code in order to guide a DSR model. Our general framework is very flexible and allows to easily swap DSR, encoder, or decoder models for the problem at hand. For diverse benchmarks and sampling conditions, MTF clearly outperforms various other training setups and models. Moreover, we found MTF to be quite robust performance-wise, and not very sensitive to precise hyper-parameter tuning.
We further highlighted on examples from neuroscience how the DSR model’s latent space exposes meaningful relations between different data modalities (neural activity and behavior in our case). In contrast to other recently proposed methods for construing a common latent space from multiple data sources (Schneider et al., 2023), our method actually delivers a generative model of the system dynamics. A crucial difference to ‘classical’ multimodal SVAEs (Kramer et al., 2022) is that our algorithm allows to optimally control trajectory and gradient flows whilst training, a key ingredient for successful DS reconstruction (Mikhaeil et al., 2022; Hess et al., 2023; Eisenmann et al., 2023). One potentially interesting future direction may be engineering ways of incorporating control signals directly at the encoder level.
@@
This way, MTF was even able, under certain conditions, to recover chaotic attractors from ordinal or categorical observations alone, including sensible estimates of their maximum Lyapunov exponent. This, to our knowledge, has never been shown before, and at first glance appears puzzling, as an ordinal or symbolic coding of the dynamics seems to remove much of the geometrical information. However, results in the field of symbolic dynamics ensure that under certain conditions a symbolic encoding of a DS preserves topological invariants and Lyapunov spectra (Osipenko, 2000; Hubler, 2012; McCullough et al., 2017; Zhang et al., 2017; Lind et al., 2021), and delay embedding theorems have been formulated for non-continuous signals like point processes (Sauer, 1994, 1995). While we foresee the major applications of MTF in truly multimodal settings, our observations here may also pave the way for interesting novel DSR applications to purely symbolic data. For instance, reconstructing dynamics from categorical consumer data or sequences of behavioral choices (on which classically reinforcement learning models would be trained) may become feasible, or even from language data (Elman, 1995), e.g. recovering the dynamics of underlying beliefs or values (Tsakalidis et al., 2022).
While here we presented first results on DS reconstruction for discrete data, whether and how much of the original state space topology of a data-generating DS can be recovered from non-continuous, non-Gaussian random variables remains an important topic for future theoretical and empirical research. Another open question remains under which conditions different modalities support (or interfere with) each other in the DSR process.
Software and Data
All code created is available at https://github.com/DurstewitzLab/MTF.
@@ Acknowledgements
This work was funded by the European Union’s Horizon 2020 Programme under grant agreement 945263 (IMMERSE), by the German Research Foundation (DFG) within the collaborative research center TRR-265 (project A06) , by the German Research Foundation (DFG) within and Germany’s Excellence Strategy EXC 2181/1 – 390900948 (STRUCTURES), and by a living lab grant by the Federal Ministry of Science, Education and Culture (MWK) of the state of Baden-Württemberg, Germany (grant number 31-7547.223-7/3/2), to DD and GK.
Impact Statement
@@ A.1 MTF framework
The MTF framework consists of four major components, 1) a DSR model for approximating the flow of the underlying, observed DS, 2) a set of modality-specific decoder models that capture the statistical particularities of the observed time series, 3) an MVAE for integrating all different data modalities in order to produce a joint latent code which can be used for TF, and 4) the actual MTF training algorithm. Here we describe each of these components in turn.
@@
While our framework is generic and works with any type of DSR model, here we used the dendPLRNN, Eq. 1, recently introduced specifically for DSR (Brenner et al., 2022) (see Appx. Table B3 for comparison with other RNN models). A major advantage of the dendPLRNN and related formulations is that it provides some degree of analytical tractability, which is of major importance in scientific settings where we are usually interested in analyzing the data-inferred model further as a formal surrogate for the real system. For example, for the dendPLRNN, we have efficient algorithms for precisely locating all fixed points, cycles, and bifurcation manifolds (Eisenmann et al., 2023), and it allows for translation into an equivalent continuous-time representation (Monfared & Durstewitz, 2020), which further eases certain types of DS analysis.
@@
To infer the latent process dendPLRNN jointly from multiple data modalities, it is connected to different decoder models that take the distinct distributional properties of each modality into account, Eqs. 2-4.
@@
Categorical observations are, like ordinal observations, not associated with a metric space, and in contrast to ordinal data, there is also no natural ordering between variable values. To couple categorical observations to the latent states, we employed the natural link function given by
@@
Here, the parameters $\bm{\beta}_{i}\in\mathbb{R}^{M\times 1}$ constitute the respective regression weights for category $i=1\ldots K-1$, with the total probability over all categories $\sum_{i=1}^{K}\pi_{i}=1$. This leads to the following log-likelihood for the categories:
$\log p_{\bm{\theta}}(\bm{O}|\bm{Z})\ =\sum_{i}^{N}\sum_{t}^{T}\sum_{k}^{K}\left[o_{it}=k\right]\log\left(\frac{\exp\left(\bm{\beta}_{k}^{T}\bm{z}t\right)}{1+\sum_{j=1}^{K-1}\exp\left(\bm{\beta}_{j}^{T}\bm{z}_{t}\right)}\right)$
where $\left[o_{it}=k\right]$ is an indicator function that is $1$ if the observation $o_{it}$ belongs to category $k$ and 00 otherwise.
For count observations $\left\{\bm{c}_{t}\right\}_{t=1}^{T}$, with $\bm{c}_{t}=\left(c_{1t},\dots,c_{Lt}\right)^{T}$, we tried three different decoder models. First, a standard Poisson model,
@@
These probabilities are related to the latent states via the log-link function, $\log\lambda_{lt}=\gamma_{0}^{(l)}+\sum_{m=1}^{M}\gamma_{m}^{(l)}z_{mt}$, where $\bm{\gamma}^{(l)}$ is a vector of coefficients. Thus, $\lambda_{lt}=e^{\gamma_{0}^{(l)}+\bm{\gamma}^{(l)}\bm{z_{t}}}$ is the expected count for the $l^{\text{th}}$ observation at time $t$. The total log-likelihood for observed counts $\bm{C}$ is then given by:
$\log p_{\bm{\theta}}(\bm{C}|\bm{Z})=\sum_{t=1}^{T}\sum_{l=1}^{L}\left[c_{lt}\log\lambda_{lt}-\lambda_{lt}-\log(c_{lt}!)\right],$ (18)
Alternatively, we tested a zero-inflated Poisson (ZIP) observation model (Lambert, 1992). The ZIP model is designed for count data that has an excess of zero counts compared to what would be expected from a traditional Poisson distribution, and thus can naturally handle overdispersion often observed in real data, such as the spike trains displayed in Fig. 5. The ZIP model addresses this by combining a binary process that determines whether a count is zero (with probability $\pi_{lt}$) with a Poisson distribution. The combined probability of an observed count given both processes is then given by
$\quad p_{\bm{\theta}}\left(c_{lt}\mid\bm{z}_{t}\right)=\left\{\begin{array}[]{cl}\pi_{lt}+\left(1-\pi_{lt}\right)e^{-\lambda_{lt}}&\text{ if }c_{lt}=0\\ \left(1-\pi_{lt}\right)\frac{\lambda_{lt}^{c_{lt}}}{c_{lt}!}e^{-\lambda_{lt}}&\text{ if }c_{lt}>0\end{array}\right.$ (19)
The probabilities are connected to the latent states via a logit and a log-link function $\log\frac{\pi_{lt}}{1-\pi_{lt}}=\beta_{0}^{(l)}+\sum_{m=1}^{M}\beta_{m}^{(l)}z_{mt}$ and $\quad\log\lambda_{lt}=\gamma_{0}^{(l)}+\sum_{m=1}^{M}\gamma_{m}^{(l)}z_{mt}$, where $\gamma^{(l)}$ and $\beta^{(l)}$are coefficient vectors. Thus,$\quad\lambda_{lt}=e^{\gamma^{(l)}z_{t}}$ is the expected count for the $l^{\text{th}}$ observation and $\pi_{lt}=\frac{e^{\beta^{(l)}z_{t}}}{1+e^{\beta^{(l)}z_{t}}}$ is the probability of observing a zero.
Finally, we tested a negative binomial model given by
@@
where $\mu_{lt}$ is the mean count and $\phi_{l}$ the dispersion parameter of the negative binomial distribution for the $l^{\text{th}}$ observation at time $t$. As for the Poisson model, we use a log-link function, $\log\mu_{lt}=\gamma_{0}^{(l)}+\sum_{m=1}^{M}\gamma_{m}^{(l)}z_{mt}$, with $\bm{\gamma}^{(l)}$ a vector of coefficients, and $z_{mt}$ the $m^{\text{th}}$ latent variable at time $t$. Properly accounting for dispersion significantly improved the modeling of the spike counts. For the evaluation we used the the likelihood function implemented in the torch.distributions.NegativeBinomial class from the PyTorch library.